abcde123:

360/15 = ________ hours 12 15 24 36

umm:

Hello, welcome to QuestionCove.

justus:

What do you think?

justus:

All you have to is divide 360 by 15 to get your answer. The word "division" means the operation which is the opposite of multiplication. In this case you have to do long division. 360 = dividend 15 = divisor Start by setting it up with the divisor 15 on the left side and the dividend 360 on the right side: 15⟌360 The divisor 15 goes into the first digit of the dividend 3, 0 times. Therefore, put 0 on top: 0 15⟌360 Multiply the divisor by the result in the previous step (15 x 0 = 0) 0 15⟌360 0 Subtract the result in the previous step from the first digit of the dividend (3 - 0 = 3) 0 15⟌360 -0 3 Move down the 2nd digit of the dividend (6): 0 15⟌360 -0 36 The divisor (15) goes into the bottom number (36), 2 time(s). Therefore, put 2 on top: 02 15⟌360 -0 36 Multiply the divisor by the result in the previous step (15 x 2 = 30): 02 15⟌360 -0 36 30 Subtract the result in the previous step from the number written above it. (36 - 30 = 6) : 02 15⟌360 -0 36 -30 6 Move down the last digit of the dividend (0): 02 15⟌360 - 0 36 - 30 60 The divisor (15) goes into the bottom number (60), 4 time(s). Therefore put 4 on top: 024 15⟌360 - 0 36 - 30 60 Multiply the divisor by the result in the previous step (15 x 4 = 60) 024 15⟌360 - 0 36 - 30 60 60 Subtract the result in the previous step from the number written above it. (60 - 60 = 0) 024 15⟌360 - 0 36 - 30 60 60 0 We are done, because there are no more digits to move down from the dividend. The answer is the top number and the remainder is the bottom number. Therefore, the answer to 360 divided by 15 calculated using Long Division is: 24